# Probing the band structure of the strongly correlated antiferromagnet NiPS3 across its phase transition

### v1

@@ Abstract
NiPS3 is an exfoliable van-der-Waals intralayer antiferromagnet with zigzag-type spin arrangement. It is distinct from other TMPS3 (TM: transition metal) materials by optical excitations into a strongly correlated state that is tied to the magnetic properties. However, the related, fundamental band structure across the antiferromagnetic phase transition has not been probed yet. Here, we use angular-resolved photoelectron spectroscopy with m resolution in combination with DFT+U calculations for that purpose. We identify a characteristic band shift across . It is attributed to bands of mixed Ni and S character related to the superexchange interaction of Ni 3t2g orbitals. Moreover, we find a structure above the valence band maximum with little angular dispersion that could not be reproduced by the calculations. The discrepancy suggests the influence of many-body interactions beyond the DFT+U approximations in striking contrast to the results on MnPS3 and FePS3, where these calculations were sufficient for an adequate description.
@@ I Introduction
Van der Waals (vdW)-type two-dimensional (2D) ferromagnets (FMs) and antiferromagnets (AFMs) Huang et al. (2017); Sun et al. (2019); Gong et al. (2017); Wang et al. (2018a); Deng et al. (2018); Fei et al. (2018) enrich the possibilities for spintronics, magneto-optical devices and magnetic sensors Li et al. (2019); Zhang et al. (2019); Burch et al. (2018); Gong and Zhang (2019); Huang et al. (2020). This is related to the fact that the 2D magnetism can be tuned via magnetic fields, electrostatic gating, current pulses, strain, light, ion intercalation, proximity and moiré lattices Burch et al. (2018); Li et al. (2019); Xu et al. (2021); Mi et al. (2022); Shin et al. (2022); Ahn et al. (2024); Cham et al. (2024) as well as due to the relatively strong magneto-electric, magneto-elastic and magneto-optic couplings Wang et al. (2022); Vaclavkova et al. (2021); Ressouche et al. (2010); Kang et al. (2020); Hwangbo et al. (2021); Dirnberger et al. (2022). Moreover, these materials enable a novel access to fundamental questions in 2D magnetism such as the Berezinskii–Kosterlitz–Thouless (BKT) transition or proximity-induced modifications of the magnetization Gong et al. (2017); Burch et al. (2018); Gong and Zhang (2019); Huang et al. (2020); Srivastava et al. (2020); Bedoya-Pinto et al. (2021); Albarakati et al. (2022); Ahn et al. (2024); Sun et al. (2024).
@@
A unique type are transition metal (TM) phosphorus trisulfides TMPS3 Joy and Vasudevan (1992); Wang et al. (2018b). In contrast to other vdW magnets, they are intralayer AFMs and, moreover, rather directly realize the three most prominent classes of 2D magnetism (Ising, XY, Heisenberg) Chittari et al. (2016); Kim et al. (2019); Autieri et al. (2022). Hence, the materials provide a versatile platform to study 2D AFM properties. The TMPS3 materials are semiconductors with a layered honeycomb structure of the TM2+ atoms, each surrounded by covalently bonded (P2S6)4- bipyramids (Fig. 1a). The magnetism arises from the competition between direct exchange interactions between neighboring TM atoms and indirect superexchange mediated by S and P atoms Autieri et al. (2022); Yan et al. (2023a); Rybak et al. (2024). Depending on the relative strength of the competitive interactions, various long-range AFM orders have been found, dubbed zigzag, Néel, and stripy AFM order Joy and Vasudevan (1992); Sivadas et al. (2015).
@@
NiPS3 stands out since it exhibits significant spin-charge coupling, leading to spin-correlated excitons whose origin is actively debated. The excitons display layer-dependent characteristics and polarization properties correlated with the material’s magnetism Kim et al. (2018); Kang et al. (2020); Lane and Zhu (2020); Wang et al. (2021); Hwangbo et al. (2021); Wang et al. (2022); Jana et al. (2023); Klaproth et al. (2023); Song et al. (2024); Wang et al. (2024); Allington et al. (2025). Notably, an ultra-sharp photoluminescence peak has been observed at approximately 1.47 eV Chittari et al. (2016); Kim et al. (2018); Muhammad et al. (2023); Du et al. (2016). Moreover, the magnetic interactions are surprisingly dominated by a third nearest neighbor AFM exchange coupling due to a substantial overlap of p- and d-orbitals along the corresponding Ni-S-P-S-Ni interaction path Scheie et al. (2023); Autieri et al. (2022); Chittari et al. (2016); Yan et al. (2023a). This leads to an intralayer AFM zigzag order as the ground state below the bulk Néel temperature 155 K, where the spins are oriented mostly in-plane along the zigzag direction (Fig. 1a) with a slight canting out-of-plane by 8∘ Wang et al. (2022). In monolayer form, the long-range magnetic order is most likely suppressed potentially leading to a BKT transition Kim et al. (2019); Sun et al. (2024). This behavior aligns with descriptions based on a highly anisotropic XXZ model Joy and Vasudevan (1992); Wildes et al. (2015).
@@
The magnetic ordering of NiPS3 has already been probed by multiple methods Kim et al. (2018, 2019); Kang et al. (2020); Hwangbo et al. (2021); Afanasiev et al. (2021); Basnet et al. (2021, 2022); Scheie et al. (2023); Muhammad et al. (2023), but none of them is sensitive to details of the band structure, which has only been probed at room temperature so far, i. e. above , using angular resolved photoelectron spectroscopy (ARPES) Yan et al. (2023b); Klaproth et al. (2023); Nitschke et al. (2025). Here, we provide band structure mapping by ARPES at variable temperature above and below leveraging the approach that we recently developed for the isostructural MnPS3 and FePS3 Strasdas et al. (2023); Pestka et al. (2024). We prevent sample charging due to photoelectrons, that often appears for semiconductors at low Bianchi et al. (2023); De Vita et al. (2022); Nitschke et al. (2023); Koitzsch et al. (2023); Voloshina et al. (2023), by exfoliating thin films of NiPS3 onto a conductive Au film deposited on Si/SiO2, itself necessary for identifying the flakes optically. The resulting band structure is compared with density functional theory calculations (DFT+U) enabling the identification of the elemental and orbital character of the different bands and revealing an optimal Hubbard parameter eV. Most importantly, we observe a pronounced band shift across for a valence band identified as a mixed Ni 3d/S 3p band that could be attributed to a next nearest neighbor superexchange path. Moreover, we reveal an additional band feature in ARPES that appears above the valence band maximum, but could not be reproduced by the DFT+U calculations. We believe that it testifies the strong correlations in NiPS3 compared to MnPS3 and FePS3 in line with the optical results Lane and Zhu (2020); Wang et al. (2022); Kim et al. (2018); Kang et al. (2020); Hwangbo et al. (2021).
@@ II Results and Discussion
NiPS3 single crystals are synthesized by the vapor transport method. Flakes are subsequently exfoliated by the scotch tape method onto Au/Ti covered Si/SiO2 at 60∘ C, shortly after the substrate was cleaned by plasma ashing Strasdas et al. (2023) (Materials and Methods). Flakes with appropriate thickness and lateral size are identified using optical and atomic force microscopy (Fig. 1b-c). Typically, these flakes exhibit various thicknesses being most thin at the rim. In this paper, we studied a 15 layer thick area with a lateral size of m2 as marked in Fig. 1b. Atomic force microscopy reveals a rms roughness of nm and no obvious contaminations in that area (Fig. 1c). The relatively small thickness is crucial to avoid charging during -ARPES at low temperature as detected, e. g. for 45 layers at 90 K, where all bands are shifted downwards by 0.3 eV.
@@
Figure 1d displays X-ray photoelectron spectroscopy (XPS) data of the 15 layer area. All components of NiPS3 appear as core level peaks Yan et al. (2021); Currò et al. (1995); Piacentini et al. (1982) as well as a S Auger peak Coad and Rivière (1972). The latter shifts with respect to the other peaks by changing the photon energy exposing the Ni 3p peak Justin Gorham (2012). A small contribution from the spin-split Au 4f peaks Justin Gorham (2012) is also visible likely originating from the substrate due to the relatively large spot of the photon beam on the sample (diameter: 15 m) at the large for XPS.
@@
The antiferromagnetic phase transition of the exfoliated NiPS3 flakes is verified by Raman spectroscopy (Fig. 1e). Due to the nearly realized D3d intralayer point group symmetry Kuo et al. (2016), NiPS3 has eight major Raman-active phonon modes, all visible in our Raman spectra with frequencies in good agreement with the literature Kim et al. (2019); Wang et al. (2019); Muhammad et al. (2023); Kuo et al. (2016). We reproduce three signatures that have previously been assigned to the AFM phase transition Kim et al. (2019); Muhammad et al. (2023). Firstly, a broad 2-magnon scattering peak, centered at 530 cm-1, appears below at 77 K and is absent at room temperature (RT) Kim et al. (2019). Secondly, the peak P2 caused by in-plane vibrations of the Ni atoms splits into a double peak (insets, Fig. 1e) as attributed to a pronounced magnon-phonon coupling Kim et al. (2019); Muhammad et al. (2023). Thirdly, the peak P8, caused by an in-plane Eg-type vibration of the (P2S6)4- bipyramids, develops a Fano-resonance line shape below T due to an interference between the phonon excitation and the continuum of 2-magnon excitations Fano (1961). These three fingerprints confirm the AFM phase transition in our NiPS3 flakes.
@@
Figure 2 shows the main result of this work, namely a band shift across T near the point. We present the raw ARPES data (a–b), the data after box-type smoothing (c, only the energy distribution curve (EDC) is shown), after additional background subtraction (d–f) and angularly averaged in -space around the point (g–h). Already the raw data (Fig. 2a–b) exhibit a slightly enhanced intensity below in between the white arrows. This change can be identified as an additional peak in the EDC at after mild smoothing (upper arrow, Fig. 2c) implying a state at eV below (: energy, : Fermi energy as probed on Au). A polynomial background subtraction (Fig. 2f) increases the visibility of that peak. The resulting energy vs. wave vector display of the photoelectron intensity (Fig. 2d–e) reveals that the peak belongs to a rather flat band extending up to an in-plane wave vector Å, where it merges with upwards dispersing bands. These features are corroborated by the angulary averaged data in space (Fig. 2g–h). The band at eV is again visible below (pink arrow, Fig. 2h). It shifts downwards by roughly 150 meV above (Fig. 2g), nearly merging with the band at eV. The insets finally show the two-dimensional plots at that energy exhibiting a ring above due to the upwards dispersing bands and a full circle below due to the additional flat band. The upwards dispersing bands visible in the main image are also slightly less steep below .
@@
Next, we compare the ARPES data with DFT+U calculations of the magnetic ground state (AFM zigzag type) and other magnetic configurations (AFM Néel and disordered). This aims to identify the elemental and orbital character of the different bands, in particular the ones that are changing across . We started by selecting the proper Hubbard parameter (DFT+U) by detailed comparison with the experimental data. The best match for the ARPES data at eV and K is found for eV (Suppl. section S2) and Å (Suppl. section S3).
@@
The band structure is additionally unfolded to the atomic Brillouin zone (grey hexagon in Fig. 3a) Popescu and Zunger (2012) (Suppl. Fig. S2). This reflects the fact that the enlarged periodicity by AFM order is barely imprinted in the wave functions Hüfner (2013). Indeed, maps of the ARPES data did not show any patterns reminiscent of the magnetic Brillouin zone, but only of the hexagonal, geometric one (Suppl. section S1). Figure 3c shows the same DFT+U data as in b, but after selecting only the orbital contributions s, pz, and d within a frame where the -direction is pointing perpendicular to the layers. According to simplified selection rules assuming plane waves as final states Moser (2017), these orbital contributions are excited preferably for our light polarization geometry Strasdas et al. (2023). Figure 3e shows the same selection of bands for the Nèel configuration (disordered magnetic configuration in Suppl. Fig. S10), while Fig. 3d shows the ARPES curvature data below . The curvature is used for better visibility of the various features (Materials and Methods section). We always cross-check with the raw data as in Fig. 2 to corroborate that a feature is not an artifact of the curvature derivation. Orbital contributions to the different bands beyond the simplified selection rules are analyzed in Suppl. section S4 with the results included in the following discussion.
@@
Multiple features are similar between ARPES (Fig. 3d) and DFT+U (Fig. 3c) as labeled by i-ix. The bright feature (i) and the nearly parabolic feature (ix) are used to align the Fermi level of the ARPES data to the DFT+U band structure. The additional structure in ARPES at eV is discussed below. The band (i) has a dominating S 3p character with contributions from Ni 3d (Supplementary figure S6). It is rather flat in calculation and experiment. The lowest energy band (ix) has a parabolic dispersion close to in both cases and a mixed P 3pz and S 3p character. This band appears quite similarly for MnPS3 and FePS3 Strasdas et al. (2023); Pestka et al. (2024). It is slightly lower in the DFT+U data than in the ARPES data. This is probably due to the facts that moves upwards with decreasing in ARPES at constant and that the band moves upwards with increasing in the DFT+U data (Suppl. Fig. S4/S5). At slightly larger are bands with nearly exclusive S 3px,y character (viii) that are barely visible in ARPES in accordance with the simplified selection rules Moser (2017). The remaining features ii-vii exhibit a mixed Ni 3d/S 3p character. In ARPES, they are dominated by a large cross-type structure gapped at (iv,vi) with a rather flat band on top (ii). A flat band of strong Ni 3d character (ii) and a gapped crossing of mixed character (iv,vi) are also visible in DFT+U. Other features are more subtle and partially distinct such as the relatively steep lower part of the cross from vi to vii in ARPES which is not continuous across the whole energy range in DFT+U. Nevertheless, the agreement in the whole energy range below eV is rather satisfactory regarding the fact that the band structure of the calculation was only rigidly shifted for adaption.
@@
This importantly implies that the magnetically induced band shift identified in Fig. 2 is at the gapped crossing of feature (iv) and (vi), more precisely it is the upper part (iv) at eV around in the ARPES data (Fig. 3d) and at eV in the DFT+U data (Fig. 3c). Consistently, these band structure features exhibit a relatively strong change in DFT+U data, if the magnetic structure is changed. The gap at is increased from eV for the zigzag configuration to 0.4 eV for the Néel configuration (Fig. 3e, box). For a disordered magnetic structure, we find that features (iv,vi) get shifted upwards by meV with respect to the zigzag configuration (Suppl. Fig. S10). Data along the direction (Suppl. Fig. S9) lead to the same conclusions.
@@
The four bands in the energy region of the observed band shift consist of all Ni 3d and S 3p orbitals as well as of a minor contribution from P 3pz (Suppl. section S4). However, only one of these bands features a bonding configuration between neighboring Ni atoms. Figure 4b shows the corresponding charge density for the state at (charge density of the other three bands, Suppl. section S5). Using the octahedral coordinate system with along a Ni-S bond, one observes Ni 3t2g orbitals (dxy, dxz, dyz) for both Ni zigzag chains, distinct by their spin orientation. On the left chain with spin , the Ni 3t2g orbitals do not have any overlap with neighboring orbitals. In contrast, for the right chain, three lobes of each Ni 3 orbital overlap with a lobe of an adjacent S 3p orbital. The zoom in Fig. 4c as well as rotated views (Suppl. movie) reveal that two of these S 3p orbitals provide a connection path between two Ni atoms (arrows in Fig. 4b), each via the two lobes of a single S 3p orbital. The third connected S 3p orbital (lower lobe of the upper Ni atom in Fig. 4c) is, however, a dead end, i. e. the S 3p orbital does not have any connection to other atoms (Suppl. movie). The final forth lobe is not connected at all and, hence, called a lone lobe as marked in Fig. 4c.
@@
The superexchange path between 2 Ni atoms by only one S orbital on both sides implies that the interaction is antiferromagnetic. This naturally explains that the band is shifted upwards in energy when these chains become ferromagnetic within the zigzag configuration making the particular bond more unfavorable. It remains unclear why we observe only this change of a band and none of the spin-polarized 3eg levels at higher energy. We conjecture that this is related to the selection rules in our ARPES geometry not being sensitive to the corresponding levels (Suppl. Fig. S6).
@@
We note in passing that we did not find any AFM induced C3 symmetry breaking in the band structure of NiPS3, neither by ARPES nor in the DFT+U data. The three different or directions are identical within error bars, albeit the antiferromagnetic zigzag structure breaks the C3 symmetry (Fig. 1a). Since the calculations do not reveal any symmetry breaking either, the effect is likely too weak to be detected. In ARPES, we can, moreover, not exclude that the relatively small single domains (1-10 m) Lee et al. (2024) partially overlap within the extraction spot of the photoelectrons (diameter: 5 m) He et al. (2024).

### v2

@@ Abstract
Abstract
NiPS3 is an exfoliable van-der-Waals intralayer antiferromagnet with zigzag-type spin arrangement. It is distinct from other TMPS3 (TM: transition metal) materials by
optical excitations into a strongly correlated state that is tied to the magnetic properties. However, the related, fundamental band structure across the antiferromagnetic phase transition has not been probed yet.
Here, we use angular-resolved photoelectron spectroscopy with m resolution in combination with DFT+U calculations for that purpose. We identify a characteristic band shift across . It is attributed to bands of mixed Ni and S character related to the superexchange interaction of Ni 3t2g orbitals. Moreover, we find a structure above the valence band maximum with little angular dispersion that could not be reproduced by the calculations. The discrepancy suggests the influence of many-body interactions beyond the DFT+U approximations in striking contrast to the results on MnPS3 and FePS3, where these calculations were sufficient for an adequate description.
@@ I Introduction
Van der Waals (vdW)-type two-dimensional (2D) ferromagnets (FMs) and antiferromagnets (AFMs) Huang2017; Sun2019b; Gong2017; Wang2018d; Deng2018; Fei2018 enrich the possibilities for spintronics, magneto-optical devices and magnetic sensors Li2019; Zhang2019; Burch2018; Gong2019; Huang2020. This is related to the fact that the 2D magnetism can be tuned via magnetic fields, electrostatic gating, current pulses, strain, light, ion intercalation, proximity and moiré lattices Burch2018; Li2019; Xu2021; Mi2022; Shin2022; Ahn2024; Cham2024 as well as due to the relatively strong magneto-electric, magneto-elastic and magneto-optic couplings Wang2022; Vaclavkova2021; Ressouche2010; Kang2020; Hwangbo2021; Dirnberger2022. Moreover, these materials enable a novel access to fundamental questions in 2D magnetism such as the Berezinskii–Kosterlitz–Thouless (BKT) transition or proximity-induced modifications of the magnetization Gong2017; Burch2018; Gong2019; Huang2020; Srivastava2020; BedoyaPinto2021; Albarakati2022; Ahn2024; Sun2024.
@@
A unique type are transition metal (TM) phosphorus trisulfides TMPS3 Joy1992; Wang2018c. In contrast to other vdW magnets, they are intralayer AFMs and, moreover, rather directly realize the three most prominent classes of 2D magnetism (Ising, XY, Heisenberg) Chittari2016; Kim2019b; Autieri2022. Hence, the materials provide a versatile platform to study 2D AFM properties. The TMPS3 materials are semiconductors with a layered honeycomb structure of the TM2+ atoms, each surrounded by covalently bonded (P2S6)4- bipyramids (Fig. 1a). The magnetism arises from the competition between direct exchange interactions between neighboring TM atoms and indirect superexchange mediated by S and P atoms Autieri2022; Yan2023; Rybak2024. Depending on the relative strength of the competitive interactions, various long-range AFM orders have been found, dubbed zigzag, Néel, and stripy AFM order Joy1992; Sivadas2015.
@@
NiPS3 stands out since it exhibits significant spin-charge coupling, leading to spin-correlated excitons whose origin is actively debated. The excitons display layer-dependent characteristics and polarization properties correlated with the material’s magnetism Kim2018; Kang2020; Lane2020; Wang2021; Hwangbo2021; Wang2022; Jana2023; Klaproth2023; Song2024; Wang2024; Allington2025. This might enable novel optomagnonic interconnects (Supplementary section S1), e. g. via the ultra-sharp exciton peak observed at approximately 1.47 eV Chittari2016; Kim2018; Muhammad2023; Du2016. It has also been shown that the properties of this exciton peak can only be described correctly by calculations employing muliplet-type exact diagonalization Klaproth2023. Moreover, the magnetic interactions are surprisingly dominated by a third nearest neighbor AFM exchange coupling due to a substantial overlap of p- and d-orbitals along the corresponding Ni-S-P-S-Ni interaction path Scheie2023; Autieri2022; Chittari2016; Yan2023. This leads to an intralayer AFM zigzag order as the ground state below the bulk Néel temperature 155 K, where the spins are oriented mostly in-plane along the zigzag direction (Fig. 1a) with a slight canting out-of-plane by 8∘ Wang2022. In monolayer form, the long-range magnetic order is most likely suppressed potentially leading to a BKT transition Kim2019b; Sun2024. This behavior aligns with descriptions based on a highly anisotropic XXZ model Joy1992; Wildes2015.
| Material | AFM
order |
Bandgap (eV) | (K) | (eV) | Band shift across |
| NiPS3
(this work) |
zigzag
(XY-like) |
1.8 | 155 | 1.6 | mixed Ni 3d/S 3p band, extra |
| band at valence band maximum | |||||
| FePS3 Pestka2024 | zigzag
(Ising-like) |
1.5 | 117 | 1.2 | S 3p-, Fe 3d- and P 3p-type bands |
| MnPS3 Strasdas2023 | Néel | 2.94 | 78 | 1.8 | Mn 3d dominated band |
@@
The magnetic ordering of NiPS3 has already been probed by multiple methods Kim2018; Kim2019b; Kang2020; Hwangbo2021; Afanasiev2021; Basnet2021; Basnet2022; Scheie2023; Muhammad2023, but none of them is sensitive to details of the band structure, which has only been probed at room temperature so far, i. e. above , using angular resolved photoelectron spectroscopy (ARPES) Yan2023b; Klaproth2023; Nitschke2025. Here, we provide band structure mapping by ARPES at variable temperature above and below leveraging the approach that we recently developed for the isostructural MnPS3 and FePS3 Strasdas2023; Pestka2024. We prevent sample charging due to photoelectrons, that often appears for semiconductors at low Bianchi2023; DeVita2022; Nitschke2023; Koitzsch2023; Voloshina2023, by exfoliating thin films of NiPS3 onto a conductive Au film deposited on Si/SiO2, itself necessary for identifying the flakes optically. The resulting band structure is compared with density functional theory calculations (DFT+U) enabling the identification of the elemental and orbital character of the different bands and revealing an optimal Hubbard parameter eV. Most importantly, we observe a pronounced band shift across for a valence band identified as a mixed Ni 3d/S 3p band that could be attributed to a next nearest neighbor superexchange path. Moreover, we reveal an additional band feature in ARPES that appears above the valence band maximum, but could not be reproduced by the DFT+U calculations. We believe that it testifies the strong correlations in NiPS3 compared to MnPS3 and FePS3 in line with the optical results Lane2020; Wang2022; Kim2018; Kang2020; Hwangbo2021.
Since the comparison between these three materials is crucial for the missing assignment of this band, we provide such a comparison in table 1. In particular, the selected values are rather similar for all three materials as expected by the identical chemical environment of the transition metal 3d levels.
@@ II Results and Discussion
NiPS3 single crystals are synthesized by the vapor transport method. Flakes are subsequently exfoliated by the scotch tape method onto Au/Ti covered Si/SiO2 at 60∘ C, shortly after the substrate was cleaned by plasma ashing Strasdas2023 (Materials and Methods). Flakes with appropriate thickness and lateral size are identified using optical and atomic force microscopy (Fig. 1b-c). Typically, these flakes exhibit various thicknesses being most thin at the rim. In this paper, we studied a 15 layer thick area with a lateral size of m2 as marked in Fig. 1b. Atomic force microscopy reveals a rms roughness of nm and no obvious contaminations in that area (Fig. 1c). The relatively small thickness is crucial to avoid charging during -ARPES at low temperature as detected, e. g. for 45 layers at 90 K, where all bands are shifted downwards by 0.3 eV.
@@
Figure 1d displays X-ray photoelectron spectroscopy (XPS) data of the 15 layer area. All components of NiPS3 appear as core level peaks Yan2021; Curr1995; Piacentini1982 as well as a S Auger peak Coad1972. The latter shifts with respect to the other peaks by changing the photon energy exposing the Ni 3p peak Gorham2012. A small contribution from the spin-split Au 4f peaks Gorham2012 is also visible likely originating from the substrate due to the relatively large spot of the photon beam on the sample (diameter: 15 m) at the large for XPS.
@@
The antiferromagnetic phase transition of the exfoliated NiPS3 flakes is verified by Raman spectroscopy (Fig. 1e). Due to the nearly realized D3d intralayer point group symmetry Kuo2016, NiPS3 has eight major Raman-active phonon modes, all visible in our Raman spectra with frequencies in good agreement with the literature Kim2019b; Wang2019; Muhammad2023; Kuo2016. We reproduce three signatures that have previously been assigned to the AFM phase transition Kim2019b; Muhammad2023. Firstly, a broad 2-magnon scattering peak, centered at 530 cm-1, appears below at 77 K and is absent at room temperature (RT) Kim2019b. Secondly, the peak P2 caused by in-plane vibrations of the Ni atoms splits into a double peak (insets, Fig. 1e) as attributed to a pronounced magnon-phonon coupling Kim2019b; Muhammad2023. Thirdly, the peak P8, caused by an in-plane Eg-type vibration of the (P2S6)4- bipyramids, develops a Fano-resonance line shape below T due to an interference between the phonon excitation and the continuum of 2-magnon excitations Fano1961. These three fingerprints confirm the AFM phase transition in our NiPS3 flakes.
@@
Figure 2 shows the main result of this work, namely a band shift across T near the point. We present the raw ARPES data (a–b), the data after box-type smoothing (c, only the energy distribution curve (EDC) is shown), after additional background subtraction (d–f) and angularly averaged in -space around the point (g–h). Already the raw data (Fig. 2a–b) exhibit a slightly enhanced intensity below in between the white arrows. This change can be identified as an additional peak in the EDC at after mild smoothing (upper arrow, Fig. 2c) implying a state at eV below (: energy, : Fermi energy as probed on Au). A polynomial background subtraction (Fig. 2f) increases the visibility of that peak. The resulting energy vs. wave vector display of the photoelectron intensity (Fig. 2d–e) reveals that the peak belongs to a rather flat band extending up to an in-plane wave vector Å, where it merges with upwards dispersing bands. These features are corroborated by the angulary averaged data in space (Fig. 2g–h). The band at eV is again visible below (Fig. 2h, pink arrow). It shifts downwards by roughly 150 meV above (Fig. 2g, green arrow), nearly merging with the band at eV (Fig. 2g-h, black arrows). The insets finally show the two-dimensional plots at that energy exhibiting a ring above due to the upwards dispersing bands and a full circle below due to the additional flat band. The upwards dispersing bands visible in the main image are also slightly less steep below .
One might suspect that the downwards shift is related to a distinct structural phase transition or to a continuous thermal contraction of the NiPS3 crystal. However, there is no structural phase transition in that temperature range as known from neutron diffraction Wildes2015, while shifts of electronic states by thermal contraction are less than 50 meV between 220 K and 45 K as deduced from electronic Raman spectroscopy Wang2022. We confirmed this by DFT+U calculations using the lattice parameters deduced from neutron diffraction at 295 K and 2 K to scale the lattice parameters in DFT+U. These calculations did not reveal any band shifts larger than 50 meV as well (not shown), in particular in the energy range where the band structure change was observed in Fig. 2.
@@
Next, we compare the ARPES data with DFT+U calculations of the magnetic ground state (AFM zigzag type) and other magnetic configurations (AFM Néel and disordered). This aims to identify the elemental and orbital character of the different bands, in particular the ones that are changing across . We started by selecting the proper Hubbard parameter (DFT+U) by detailed comparison with the experimental data. The best match for the ARPES data at eV and K is found for eV (Suppl. section S3) and Å (Suppl. section S4). The selection is corroborated by photon energy dependent ARPES data that are successfully compared with the dependent DFT+U data employing an inner potential of 12.1 eV (Suppl. Fig. S5).
@@
The band structure is additionally unfolded to the atomic Brillouin zone (grey hexagon in Fig. 3a) Popescu2012 (Suppl. Fig. S2). This reflects the fact that the enlarged periodicity by AFM order is barely imprinted in the wave functions Hüfner2013. Indeed, maps of the ARPES data did not show any patterns reminiscent of the magnetic Brillouin zone, but only of the hexagonal, geometric one (Suppl. section S2). Figure 3c shows the same DFT+U data as in b, but after selecting only the orbital contributions s, pz, and d within a frame where the -direction is pointing perpendicular to the layers. According to simplified selection rules assuming plane waves as final states Moser2017, these orbital contributions are excited preferably for our light polarization geometry Strasdas2023. Orbital contributions to the different bands beyond the simplified selection rules are analyzed in Suppl. section S5 with the results partially included in the following discussion. Figure 3e shows the same selection of bands for the Nèel configuration which is not realized in NiPS3. This additional band structure serves to identify bands that are changing with the magnetic arrangement of the atoms. For the same purpose, we also provide the calculated band structure of a disordered magnetic configuration in Suppl. Fig. S11. Finally, Fig. 3d shows the ARPES curvature data below . The curvature is used for better visibility of the various features (Suppl. section S12). We always cross-check with the raw data as in Fig. 2 to corroborate that a feature is not an artifact of the curvature derivation.
@@
Multiple features are similar between the ARPES results (Fig. 3d) and the DFT+U data of the zigzag configuration (Fig. 3c) as labeled by i-ix. The bright feature (i) and the nearly parabolic feature (ix) are used to align the Fermi level of the ARPES data to the DFT+U band structure. The additional structure in ARPES at eV is discussed below. The band (i) has a dominating S 3p character with contributions from Ni 3d (Suppl. Fig. S6). It is rather flat in calculation and experiment. The lowest energy band (ix) has a parabolic dispersion close to in both cases and a mixed P 3pz and S 3p character. This band appears quite similarly for MnPS3 and FePS3 Strasdas2023; Pestka2024. It is slightly lower in the DFT+U data than in the ARPES data. This is probably due to the facts that firstly moves upwards with decreasing in ARPES at constant and secondly this band moves upwards with increasing in the DFT+U data (Suppl. Figs. S4/S5). At slightly larger are bands with nearly exclusive S 3px,y character (viii) that are barely visible in ARPES in accordance with the simplified selection rules Moser2017. The remaining features ii-vii exhibit a mixed Ni 3d/S 3p character. In ARPES, they are dominated by a large cross-type structure gapped at (iv,vi) with a rather flat band on top (ii). A flat band of strong Ni 3d character (ii) and a gapped crossing of mixed character (iv,vi) are also visible in DFT+U. Other features are more subtle and partially distinct such as the relatively steep lower part of the cross from vi to vii in ARPES which is not continuous across the whole energy range in DFT+U. Nevertheless, the agreement in the whole energy range below eV is rather satisfactory regarding the fact that the band structure of the calculation was only rigidly shifted for adaption.
@@
This importantly implies that the magnetically induced band shift identified in Fig. 2 is at the gapped crossing of feature (iv) and (vi), more precisely it is the upper part (iv) at eV around in the ARPES data (Fig. 3d). Accordingly, it is at eV in the DFT+U data (arrow in Fig. 3c). Consistently, these band structure features exhibit a relatively strong change in DFT+U data, if the magnetic structure is changed. For the zigzag configuration, the gapped area at is eV wide (arrow in Fig. 3c), while it is increased to eV for the Néel configuration (Fig. 3e, within dotted box). For a disordered magnetic structure, we find that features (iv,vi) get shifted upwards by meV with respect to the zigzag configuration (Suppl. Fig. S11). Data along the direction are very similar to the direction (Suppl. Fig. S9) and, hence, support all our conclusions.
@@
The four bands in the energy region of the observed band shift consist of all Ni 3d and S 3p orbitals as well as of a minor contribution from P 3pz (Suppl. section S5). However, only one of these bands features a bonding configuration between neighboring Ni atoms. Figure 4b shows the corresponding charge density for the state at (charge density of the other three bands, Suppl. section S6). Using the octahedral coordinate system with along a Ni-S bond, one observes Ni 3t2g orbitals (dxy, dxz, dyz) for both Ni zigzag chains, distinct by their spin orientation. On the left chain with spin , the Ni 3t2g orbitals do not have any overlap with neighboring orbitals. In contrast, for the right chain, three lobes of each Ni 3t2g orbital overlap with a lobe of an adjacent S 3p orbital. The zoom in Fig. 4c as well as rotated views (Suppl. movie) reveal that two of these S 3p orbitals provide a connection path between two Ni atoms (arrows in Fig. 4b), each via the two lobes of a single S 3p orbital. The third connected S 3p orbital (lower lobe of the upper Ni atom in Fig. 4c) is, however, a dead end, i. e. the S 3p orbital does not have any connection to other atoms (Suppl. movie). The final forth lobe of the Ni 3t2g orbital is not connected at all and, hence, called a lone lobe as marked in Fig. 4c.
@@
The superexchange path between two Ni atoms by only one S orbital on both sides implies that the interaction is antiferromagnetic. This naturally explains that the band is shifted upwards in energy when these chains become ferromagnetic within the zigzag configuration making the particular bond more unfavorable. It remains unclear why we observe only this change of a band and none of the spin-polarized 3eg levels at higher energy. We conjecture that this is related to the selection rules in our ARPES geometry not being sensitive to the corresponding levels (Suppl. Fig. S6).
@@
We note in passing that we did not find any AFM induced C3 symmetry breaking in the band structure of NiPS3 by ARPES. The three different or directions are identical within error bars, albeit the antiferromagnetic zigzag structure breaks the C3 symmetry (Fig. 1a). Since the DFT+U calculations do not reveal any symmetry breaking that is significantly larger than the energy resolution of the experiment (50 meV), the effect is likely too weak to be detected (Suppl. section S8). In ARPES, we can, moreover, not exclude that the relatively small single domains (1-10 m) Lee2024 partially overlap within the extraction spot of the photoelectrons (diameter: 5 m) He2024.